# Graph Chirp Signal and Graph Fractional Vertex-Frequency Energy Distribution

### v1

@@ Abstract
Graph signal processing (GSP) has emerged as a powerful framework for analyzing data on irregular domains. In recent years, many classical techniques in signal processing (SP) have been successfully extended to GSP. Among them, chirp signals play a crucial role in various SP applications. However, graph chirp signals have not been formally defined despite their importance. Here, we define graph chirp signals and establish a comprehensive theoretical framework for their analysis. We propose the graph fractional vertex-frequency energy distribution (GFED), which provides a powerful tool for processing and analyzing graph chirp signals. We introduce the general fractional graph distribution (GFGD), a generalized vertex-frequency distribution, and the reduced interference GFED, which can suppress cross-term interference and enhance signal clarity. Furthermore, we propose a novel method for detecting graph signals through GFED domain filtering, facilitating robust detection and analysis of graph chirp signals in noisy environments. Moreover, this method can be applied to real-world data for denoising more effective than some state-of-the-arts, further demonstrating its practical significance.
@@
Chirp signal, filtering, general fractional graph distribution, graph fractional vertex-frequency energy distribution, signal detection.
@@ I Introduction
Graph signal processing (GSP) is an emerging field that analyzes and processes signals defined on irregular graph structures [1, 2, 3]. It extends traditional signal processing (SP), which typically operates in Euclidean domains, to non-Euclidean domains, facilitating the analysis of data from networks such as social, sensor, and biological systems, where traditional SP may be inadequate [1, 4, 5, 6, 7].
@@
In SP, the fractional Fourier transform (FRFT) [8, 9] is a generalization of the classical Fourier transform (FT) [10, 11] that facilitates a more flexible analysis of signals in different fractional domains. The FRFT is an effective tool for analyzing and processing linear frequency modulation (LFM) signals, also known as chirp signals, which are characterized by their frequency variation over time [12, 13, 8, 14, 15]. In GSP, the graph Fourier transform (GFT) [5, 4, 2, 16] can be extended to graph fractional Fourier transform (GFRFT) [17, 18] by introducing the fractional parameter $a$, thereby facilitating a more sophisticated signal analysis in fractional domains defined on graphs. While the FRFT has been extensively studied for chirp signals, the concept of graph chirp signal remains absent in the realm of GSP. Defining such a signal is crucial for enhancing the analytical capabilities of GSP.
@@
Time-frequency analysis is a fundamental technique in SP [19, 20]. Time-frequency distributions provide simultaneous representations of signals’ attributes in the combined time-frequency domain, enabling a more comprehensive analysis of signals with time-varying frequency content. Among these distributions, the Wigner-Ville distribution (WD) stands out as one of the most widely used [21, 22, 23], due to its high resolution, making it particularly effective for capturing the instantaneous frequency and temporal dynamics of signals. In the realm of GSP, however, there is no direct extension of the WD for analyzing signals defined on graphs. Despite this, the graph vertex-frequency energy distribution (GED) can fulfill a similar role. In GSP, the GED [24, 25], known as the vertex-frequency distribution, characterizes the energy distribution within a graph signal across its vertices and frequency components. This distribution is analogous to the Rihaczek distribution [26, 27] in time-frequency analysis.
In classical SP, the relationship between WD and FRFT reflects a rotation in the time-frequency plane [8], making it particularly useful for analyzing chirp signals. Fractionalizing the WD enriches the structure of the time-frequency plane, thereby enhancing the performance of signal analysis, such as in signal detection and filtering [28, 29, 30, 31, 32]. Conversely, no analogous rotational relationship exists between the GED and the GFRFT. Thus, these methods are not typically suitable for analyzing graph chirp signals. However, fractionalizing the GED enhances the structure of the vertex-frequency plane, making it more complex and enriched, and thus better suited for analyzing graph chirp signals.
In SP, the Cohen’s class time-frequency distribution (CD) is one of the most comprehensive bi-linear kernel function time-frequency distributions in SP [20]. When the kernel function is set to 1, the CD degenerates into the well-known WD, which, despite its widespread applications, still contains cross-terms that can introduce interference in time-frequency plane. Thus, researchers have focused on designing different kernel functions to suppress these cross-terms [33, 34, 35, 36, 37, 38]. The CD in GSP aligns with the general graph distribution (GGD). Likewise, the GGD can be fractionalized, facilitating a more nuanced representation in the vertex-frequency domain. This fractionalization helps to better manage and suppress cross-terms interference, improving the accuracy of graph signal analysis.
@@
In summary, the analysis of graph chirp signals is crucial in GSP. The main objective of this paper is to define graph chirp signals and provide a tool for processing them by extending the GED and integrating it with the GFRFT to develop a fractionalized GED, thereby enhancing the theoretical framework for graph signal analysis. This approach offers an effective framework for analyzing and processing graph chirp signals with the potential to significantly impact the field of SP and graph theory, underscoring its significance to the audience. The main contributions of this paper are summarized as follows:
  • This paper defines the graph chirp signal.
  • This paper defines the graph fractional vertex-frequency energy distribution (GFED) and presents its corresponding properties.
  • This paper defines the general fractional graph distribution (GFGD) and presents its corresponding properties. Moreover, this paper gives some examples of reduced interference GFED.
  • This paper proposes a method for chirp signal detection using filtering in the GFED domain. Furthermore, the method can be applied to real-world scenarios for denoising more effective than some state-of-the-arts.
@@
The remainder of this paper is organized as follows. Section II introduces some preliminary concepts. Section III defines graph chirp signals and discusses their relevant properties. Section IV presents the definition and properties of the GFED. Section V introduces the GFGD and the reduced interference GFED. Section VI proposes a method for detecting graph chirp signals using filtering in the GFED domain. Section VII provides numerical experiments on the detection of graph chirp signals and the filtering of real-world data. Finally, Section VIII concludes the paper. Fig. 1 presents a mind map delineating the framework and principal concepts of this paper. All the technical proofs of our theoretical results are relegated to the Appendix parts.
@@ II-A Graph Fourier transform
Let $\mathcal{G}=\left\{\mathcal{N},\mathcal{E},\mathbf{A}\right\}$ be a graph or network with a set of $N$ nodes $\mathcal{N}$, edges $\mathcal{E}$ such that $(m,n)\in\mathcal{E}$ and the graph adjacency $\mathbf{A}$. If there is no edge from node $n$ to node $m$, the element $\mathbf{A}_{mn}=0$, otherwise $\mathbf{A}_{mn}=1$. The matrix $\mathbf{W}$ represents the strength of the connections between the nodes in the graph, where each element $W_{mn}$ denotes the weight of the edge between nodes $m$ and $n$. The diagonal elements $W_{nn}$ are zeros. The Laplacian is defined as $\mathbf{L=D-W}$, where $\mathbf{D}$ is the diagonal degree matrix with $D_{nn}=\sum_{m=1}^{N}W_{nm}$. For any undirected graph, the Laplacian is symmetric, i.e. $\mathbf{L}=\mathbf{L}^{\rm T}$. A graph signal $\mathbf{x}=[x(1),x(2),\cdots,x(N)]^{\mathrm{T}}$, $\mathbf{x}\in\mathbb{C}^{N}$, is defined as a mapping form the set of vertices $\mathcal{N}$ to $\mathbb{C}^{N}$, i.e. $\mathcal{N}_{n}\rightarrow x(n)$.
In [2], the GFT is defined using the general graph shift operator (GSO) $\mathbf{Z}$ (e.g. adjacency $\mathbf{A}$, weighted adjacency $\mathbf{W}$, Laplacian $\mathbf{L}$, row normalized adjacency, summetric normalized Laplacian). The GSO $\mathbf{Z}$ can be decomposed as $\mathbf{Z=U\Lambda U}^{-1}$, where $\mathbf{U}$ is the matrix whose columns are the eigenvectors $\mathbf{u}_{k}$, $k=1,2,\cdots,N$, and $\mathbf{\Lambda}$ is the diagonal matrix consisting of eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{N}$ on the diagonal. It is obvious that $\mathbf{U}$ is unitary. For clarity, unless explicitly stated, the decomposition is based on the graph Laplacian in this paper.
@@ II-B Graph fractional Fourier transform
$\mathbf{F}_{G}^{a}=\mathbf{P}\mathbf{J}_{F}^{a}\mathbf{P}^{-1}$ (3)
@@ III The graph chirp signal
LFM signal, also known as chirp signal, characterized by a frequency that varies with time, have long been utilized in classical SP for applications such as radar, communications, and system identification. A chirp signal typically exhibits a frequency that changes linearly with time.
In the continuous domain, a LFM signal is defined as [12, 13]
$f(t)={\rm e}^{{\rm i}(f_{0}t+\frac{f_{k}}{2}t^{2})}$ (6)
where $f_{0}$ and $f_{k}$ denote the initial frequency and chirp rate, respectively.
As is known to all, the fractional Fourier transform [39, 40] of a chirp signal generates a characteristic spike, often referred to as an impulse. Thus, we can extend the idea of chirp signals to the graph domain and define the graph chirp signal.
@@
Definition 1: For a graph or network $\mathcal{G}$ with the GFT matrix $\mathbf{F}_{G}$, the graph chirp signal $\mathbf{u}_{k}^{a}$ is defined as the $k$th column of the inverse of $a$ order GFRFT matrix, as follows:
$\mathbf{u}_{k}^{a}=\left(\mathbf{F}_{G}^{a}\right)^{-1}(:,k),$ (7)
@@
where $a$ is the graph chirp rate and $k$ is the graph initial frequency.
To provide a clearer illustration of the behavior of chirp signals, in Fig. 2, we give examples on four different graph structures: the David sensor, Swiss roll, comet, and low-stretch tree. For each graph, we examine the behavior of chirp signals with varying parameters, specifically the chirp rates $a=\{0.3,0.5,0.8\}$, as well as initial frequencies $k=\{1,30,50\}$. Each graph contains $N=64$ vertices.
$k=1$
$a=0.3$
$k=30$
$k=50$
$k=1$
$k=30$
$k=50$
$a=0.5$
$a=0.8$
(a) David sensor
(b) Swiss roll
$a=0.3$
$a=0.5$
$a=0.8$
(c) Comet
(d) Low stretch tree
Figure 2: Graph chirp signals with different chirp rates and initial frequencies on four graph structers. (a) David sensor. (b) Swiss roll. (c) Comet. (d) Low stretch tree.
@@
$\mathbf{A}=\left[\begin{array}[]{llll}1&&&1\\ &\ddots&&\\ &&1&\end{array}\right].$ (11)
On a cycle graph, the GFRFT matrix, $\mathbf{F}_{G}^{a}$, is equal to the discrete fractional Fourier transform (DFRDT) matrix [17]. Consequently, the graph chirp signals on a cycle graph are defined as the columns of the DFRFT matrix with a fractional order $-a$. Fig. 3 illustrates the graph chirp signals on a cycle graph with $N=20$ vertices for chirp rates $a=\{0.3,0.5,0.8\}$ and initial frequencies $k=\{1,9,18\}$.
$k=1$
$a=0.3$
$k=9$
$k=18$
$a=0.5$
$a=0.8$
Figure 3: Graph chirp signals with different chirp rates and initial frequencies on cycle graph.
The GFRFT of a graph chirp signal $\mathbf{u}_{k}^{a}$ produces a corresponding impulse in the graph fractional spectral domain, analogous to the chirp signal’s behavior in classical SP. For simplicity, let $\mathbf{U}_{a}=(\mathbf{u}_{1}^{a},\mathbf{u}_{2}^{a},\cdots,\mathbf{u}_{N}^{a})$.
For example, Fig. 4 illustrates the GFRFT of the chirp signal $\mathbf{u}_{50}^{0.5}$ on the David sensor, with the fractional order $a$ varying from $0.1$ to $1$. As can be seen, the GFRFT for this fractional order generates an impulse at the chirp rate $a=0.5$, indicating that the signal exhibits a well-defined frequency concentration at this order. At other fractional orders, the spectrum is more mixed. The GFRFT effectively handles graph chirp signals by providing a flexible method to adjust the order. This aligns with the classical treatment of chirp signals, where the frequency evolution over time is captured precisely.
Figure 4: GFRFTs of the chirp signal on David sensor ($a=0.5$, $k=50$) at different fractional orders.
IV Fractional vertex-frequency energy distribution
IV-A Vertex-frequency energy distribution
In classical SP, the WD [21, 22] is a well-known tool that provides a joint representation of a signal, offering valuable insight into its time-frequency characteristics. It is particularly effective for the analysis and processing of LFM signals, such as filtering and detection. In GSP, the concept of energy distribution holds an analogous importance to the role of WD in classical SP.
In [24], the GED is proposed, which corresponds to the Rihaczek distribution in time-frequency analysis and is defined as follows:

### v2

# Graph Chirp Signal and Graph Fractional Vertex–Frequency Energy Distribution

@@ Abstract
Graph signal processing (GSP) has emerged as a powerful framework for analyzing data on irregular domains. In recent years, many classical techniques in signal processing (SP) have been successfully extended to GSP. Among them, chirp signals play a crucial role in various SP applications. However, graph chirp signals have not been formally defined despite their importance. Here, we define graph chirp signals and establish a comprehensive theoretical framework for their analysis. We propose the graph fractional vertex–frequency energy distribution (GFED), which provides a powerful tool for processing and analyzing graph chirp signals. We introduce the general fractional graph distribution (GFGD), a generalized vertex–frequency distribution, and the reduced interference GFED, which can suppress cross-term interference and enhance signal clarity. Furthermore, we propose a novel method for detecting graph signals through GFED domain filtering, facilitating robust detection and analysis of graph chirp signals in noisy environments. Moreover, this method can be applied to real-world data for denoising more effective than some state-of-the-arts, further demonstrating its practical significance.
@@
Chirp signal, filtering, general fractional graph distribution, graph fractional vertex–frequency energy distribution, signal detection.
@@ I Introduction
Graph signal processing (GSP) is an emerging field that analyzes and processes signals defined on irregular graph structures [1, 2, 3]. It extends traditional signal processing (SP), which typically operates in Euclidean domains, to non–Euclidean domains, facilitating the analysis of data from networks such as social, sensor, and biological systems, where traditional SP may be inadequate [1, 4, 5, 6, 7].
@@
In SP, the fractional Fourier transform (FRFT) [8, 9] is a generalization of the classical Fourier transform (FT) [10, 11] that facilitates a more flexible analysis of signals in different fractional frequency domains. The FRFT is an effective tool for analyzing and processing linear frequency modulation (LFM) signals, also known as chirp signals, which are characterized by their frequency variation over time [12, 13, 8, 14, 15]. In GSP, the graph Fourier transform (GFT) [5, 4, 2, 16] has been extended to graph fractional Fourier transform (GFRFT) [17, 18] by introducing the fractional parameter $a$, facilitating a more sophisticated signal analysis in fractional graph spectral domains. However, the class of graph signals particularly suited for analysis using GFRFT remains unexplored in the current GSP literature. Motivated by the close connection between FRFT and chirp signals, we can define the graph chirp signals as natural counterparts in the graph domain.
@@
Time–frequency analysis is a fundamental technique in SP [19, 20]. Time–frequency distributions provide simultaneous representations of signals’ attributes in the combined time–frequency domain, enabling a more comprehensive analysis of signals with time-varying frequency content. Among these, the Wigner-Ville distribution (WD) stands out as one of the most widely used methods [21, 22, 23], owing to its high resolution and capability to accurately capture instantaneous frequency and temporal dynamics. The relationship between WD and FRFT reflects a rotation of the time–frequency plane [8], making it particularly useful for analyzing chirp signals.
However, as explicitly pointed out by Stankovic and Sejdic in [24], the WD is not directly suitable for extension to GSP framework. Within GSP, the graph vertex–frequency energy distribution (GED) [25, 26] has emerged as a more appropriate alternative, which serves as the graph counterpart of the classical Rihaczek distribution [27, 28]. The GED characterizes how the energy of a graph signal is distributed jointly over the vertex and frequency domains. Nevertheless, unlike the classical scenario, there is no corresponding rotational relationship between GED and GFRFT, primarily due to the non–Euclidean nature of graph domains and the generally non–idempotent property of the GFT matrix. As a result, the GED is not well-suited for analyzing graph chirp signals. To address this challenge, we introduce the graph fractional vertex–frequency energy distribution (GFED) by extending the GFT operator in GED to its fractional counterpart, the GFRFT. The GFED can capture the vertex–fractional-frequency representations, making it particularly suitable for nuanced analysis of graph chirp signals. Analogous to the fractional Rihaczek distribution in SP, the GFED can be regarded as a natural extension from traditional SP to GSP.
When dealing with multi–component chirp signals, both the Rihaczek distribution and its fractional counterpart suffer from cross-term interference. These cross-terms not only obscure the true signal components but also significantly degrade the time–frequency resolution, hindering accurate interpretation. To address this issue, the Cohen’s class time–frequency distribution (CD) [20] is developed as a comprehensive framework for constructing bi–linear kernel–based representations, with the Rihaczek distribution being one of its specific instances. Within this framework, extensive research has focused on designing adaptive kernel functions to effectively suppress cross-terms [29, 30, 31, 32, 33, 34]. The graph domain analogue of the CD is the general graph distribution (GGD). However, the GGD has similar limitations with GED that it struggles to effectively represent multi-component graph chirp signals. To overcome this limitation, we can replace the GFT operator within GGD with the GFRFT operator, obtaining the general fractional graph distribution (GFGD). This new distribution enhances the representation of graph chirp signals and improves cross-term suppression. Hence, the GFGD not only generalizes the fractional CD [35, 36, 37] to vertex–fractional-frequency domains in GSP but also constitutes a substantive extension of the GGD, tailored to address the cross-term interference of graph chirp signals.
@@
In summary, the analysis of graph chirp signals is crucial in GSP. The main objective of this paper is to define graph chirp signals and provide a tool for processing them by extending the GED and integrating it with the GFRFT to develop a fractionalized GED, thereby enhancing the theoretical framework for graph signal analysis. This approach offers an effective framework for analyzing and processing graph chirp signals with the potential to significantly impact the field of SP and graph theory. The main contributions of this paper are summarized as follows:
  • Define the graph chirp signal and derive its properties.
  • Define the GFED and derive its properties.
  • Define the GFGD, derive its properties, and provide some examples of reduced interference GFED.
  • Develop a GFED-based chirp detection method via filtering, apply it to real-world denoising tasks, and achieve superior performance to some state-of-the-art methods.
@@
The remainder of this paper is organized as follows. Section II introduces some preliminary concepts. Section III defines graph chirp signals and discusses their relevant properties. Section IV presents the definition and properties of the GFED. Section V introduces the GFGD and the reduced interference GFED. Section VI proposes a method for detecting graph chirp signals using filtering in the GFED domain. Section VII provides numerical experiments on the detection of graph chirp signals and the filtering of real-world data. Section VIII discusses modeling real–world signals with graph chirp signals and outlines possible extension. Finally, Section IX concludes the paper. Fig. 1 presents a mind map delineating the framework and principal concepts of this paper. All the technical proofs of our theoretical results are relegated to the Appendix parts.
@@ II-A Graph Fourier transform
Let $\mathcal{G}=\left\{\mathcal{N},\mathcal{E},\mathbf{A}\right\}$ be a graph or network with a set of $N$ nodes $\mathcal{N}$, edges $\mathcal{E}$ such that $(m,n)\in\mathcal{E}$ and the graph adjacency $\mathbf{A}$. If there is no edge from node $n$ to node $m$, the element $\mathbf{A}_{mn}=0$, otherwise $\mathbf{A}_{mn}=1$. The matrix $\mathbf{W}$ represents the strength of the connections between the nodes in the graph, where each element $W_{mn}$ denotes the weight of the edge between nodes $m$ and $n$. The diagonal elements $W_{nn}$ are zeros. The Laplacian is defined as $\mathbf{L=D-W}$, where $\mathbf{D}$ is the diagonal degree matrix with $D_{nn}=\sum_{m=1}^{N}W_{nm}$. For any undirected graph, the Laplacian is symmetric, i.e. $\mathbf{L}=\mathbf{L}^{\rm T}$. A graph signal $\mathbf{x}=[x(1),x(2),\cdots,x(N)]^{\mathrm{T}}$, $\mathbf{x}\in\mathbb{C}^{N}$, is defined as a mapping from the set of vertices $\mathcal{N}$ to $\mathbb{C}^{N}$, i.e. $\mathcal{N}_{n}\rightarrow x(n)$.
In [2], the GFT is defined using the general graph shift operator (GSO) $\mathbf{Z}$ (e.g. adjacency $\mathbf{A}$, weighted adjacency $\mathbf{W}$, Laplacian $\mathbf{L}$, row normalized adjacency, summetric normalized Laplacian). In this paper, we considered the case where the underlying graph is undirected, so that the GSO $\mathbf{Z}$ is symmetric and diagonalizable. In particular, it can be decomposed as $\mathbf{Z}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{-1}$. Here, $\mathbf{U}$ is the matrix whose columns are the eigenvectors $\mathbf{u}_{k}$, $k=1,2,\cdots,N$, and $\mathbf{\Lambda}$ is the diagonal matrix consisting of eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{N}$ on the diagonal. It is obvious that $\mathbf{U}$ is unitary. For clarity, unless explicitly stated, the decomposition is based on the graph Laplacian in this paper.
@@ II-B Graph fractional Fourier transform
$\mathbf{F}_{G}^{a}=\mathbf{P}\mathbf{J}_{F}^{a}\mathbf{P}^{-1}.$ (3)
@@ III The graph chirp signal
LFM signal, also known as chirp signal, characterized by a frequency that varies with time, have long been utilized in classical SP for applications such as radar, communications, and system identification [38, 15, 39]. A chirp signal typically exhibits a frequency that changes linearly with time.
In the continuous domain, a chirp signal is defined as [12, 13]
$f_{f_{0},f_{k}}(t)={\rm e}^{{\rm i}(f_{0}t+\frac{f_{k}}{2}t^{2})},$ (6)
where $f_{0}$ and $f_{k}\neq 0$ denote the initial frequency and chirp rate, respectively.
We list several fundamental properties of chirp signals below.
Property 1 (FRFT–invariance) Applying the FRFT to a chirp signal results in another chirp signal [40, 41]. When the chirp rate $f_{k}$ and the FRFT angle $\alpha$ satisfy the condition $f_{k}=-\mathrm{cot}\alpha$, the FRFT of the chirp signal becomes an impulse function at the frequency $w=\frac{f_{0}}{\mathrm{csc}^{2}\alpha}$.
Property 2 (Constant amplitude) Chirp signals have constant instantaneous amplitude. For all time $t\in\mathbb{R}$,
$\left|f_{f_{0},f_{k}}(t)\right|^{2}=1.$ (7)
Property 3 (Orthogonality) Chirp signals with different initial frequencies are orthogonal, , satisfying
$\langle f_{f_{0},f_{k}},f_{f_{1},f_{k}}\rangle=2\pi\delta_{f_{0},f_{1}},$ (8)
[…]
Property 4 (Time–frequency domain energy concentration) The fractional WD with order $\alpha$ [40, 41] of a chirp signal $f_{f_{0},f_{k}}(t)$ is given by
[…]
$=\sqrt{\rm i\,tan\alpha+1}\,\mathrm{e}^{\mathrm{i}\frac{1}{2}w^{2}\mathrm{cot}\alpha}\mathrm{e}^{-\mathrm{i}\frac{\left(f_{0}+f_{k}t-w\mathrm{csc}\alpha\right)}{2\mathrm{cot}\alpha}},$ (9)
where $K_{\alpha}(\tau,w)=\sqrt{\frac{1-\mathrm{i\,cot}\alpha}{2\pi}}\mathrm{e}^{\mathrm{i}\frac{1}{2}\left(\tau^{2}+w^{2}\right)\mathrm{cot}\alpha-\mathrm{i}w\tau{\rm csc}\alpha}$ is the fractional kernel function. In particular, when ${\rm cot}\alpha=0$, it reduces to the standard WD
$W_{f}(t,w)=\sqrt{2\pi\left(1-\mathrm{i\,cot}\alpha\right)}\,\delta\left(w-(f_{0}+f_{k}t)\right),$ (10)
which shows that the energy of a chirp signal lies along a line $w=f_{0}+f_{k}t$ in the time–frequency plane.
In SP, chirp signals are characterized by their instantaneous frequency being a linear function of time. However, this concept can not naturally extend to graph domains, where the vertex set is typically unordered and non–Euclidean, making it infeasible to define instantaneous frequency as a linear function of vertex index. Thus, a direct extension of the classical definition is not applicable.
Fortunately, as shown in Property 1, chirp signals also possess a critical inherent feature: they are the inverse FRFT of a delta function when the fractional order matches the chirp rate. Due to the reversibility of the FRFT, this means that chirp signals are uniquely characterized by the fact that their FRFT is a delta function, and vice versa. This feature is intrinsic to chirp signals.
Motivated by this fundamental insight and the properties of classical chirp signals, we can extend the idea of chirp signals to the graph domain and define the graph chirp signal.
@@
Definition 1: For a graph or network $\mathcal{G}$ with the GFT matrix $\mathbf{F}_{G}$, the graph chirp signal $\mathbf{u}_{k}^{a}$ ($a\neq 0$) is defined as the inverse GFRFT with order $a$ of $\mathbf{e}_{k}$ as follows:
$\mathbf{u}_{k}^{a}=\left(\mathbf{F}_{G}^{a}\right)^{-1}\mathbf{e}_{k},$ (11)
@@
where $a$ is the graph chirp rate, $k$ is the graph initial frequency and $\mathbf{e}_{k}$ denotes the standard basis vector with 1 at the $k$-th position and 0 elsewhere, i.e. $[\mathbf{e}_{k}]_{i}=\delta_{ik}$.
Theorem 1 (GFRFT–invariance): The graph chirp signal $\mathbf{u}_{k}^{a}$ remains a graph chirp under the GFRFT with order $b$, i.e.,
$\widehat{\left(\mathbf{u}_{k}^{a}\right)}_{b}=\mathbf{u}_{k}^{a-b}.$ (12)
In particular, when $a=b$, we have
$\widehat{\left(\mathbf{u}_{k}^{a}\right)}_{a}=\mathbf{e}_{k}.$ (13)
Proof: See Appendix A.
The GFRFT–invariance corresponds to the Property 1 of chirp signals. This stability under the GFRFT can also be interpreted as chirp rate additivity, that is
$\mathbf{u}_{k}^{a+b}=\mathbf{F}_{G}^{-b}\mathbf{u}_{k}^{a}.$ (14)
Theorem 2 (Constant–norm): The graph chirp signals are unit-norm in the $\ell_{2}$-sense
$\|\mathbf{u}_{k}^{a}\|_{2}=1.$ (15)
Proof: See Appendix A.
This implies that the energy of graph chirp signals is invariant with respect to the chirp rates and graph initial frequencies. Furthermore, this property complements Property 1 that since the GFRFT of order $b$ maps a graph chirp of rate $a$ to another graph chirp of rate $a-b$ and all such graph chirp signals are unit-norm, the total signal energy is preserved under successive GFRFT operations. Although GSP is inherently defined over discrete graph structures, the graph chirp rate is a continuous parameter. As a result, varying $a$ and $b$ continuously gives rise to a dense family of graph chirp signals and their transformations. Therefore, the preservation of energy across these transformations can be interpreted as an approximate continuity under GFRFT operations, despite the discrete nature of the underlying graph structure. The constant-norm of graph signals correspond to the Property 2 of chirp signals.
Theorem 3 (Orthogonality): The graph chirp signals associated with different graph initial frequencies are orthogonal when the GFRFT matrix is unitary
$\langle\mathbf{u}_{k}^{a},\mathbf{u}_{l}^{a}\rangle=\delta_{kl}.$ (16)
Proof: See Appendix A.
It is obvious that when $k=l$, this identity reduces to the unit–norm condition established in Theorem III. Due to the orthogonality of the GFRFT basis, the set $\left\{\mathbf{u}_{k}^{a},k=1,2,\cdots,N\right\}$ forms an orthonormal basis, enabling graph signals to be efficiently decomposed, projected, and analyzed in the graph spectral domain. This result parallels the Property 3 of chirp signals, which asserts orthogonality between chirp signals of different initial frequencies.
Comparisons with the graph chirp signals defined in [42]: A related concept of graph chirp signal was previously introduced by Chen et al. in their work [42]. Rather than generalizing the definition in time domain directly — which is infeasible due to the lack of a natural vertex ordering — they define the graph chirp signals from the perspective of graph spectral domain. Their method emulates the spectral form in the GFT domain of classical chirp signals, with the objective of approximating the time domain expression described in (6). In contrast, our approach originates from the inherent properties of chirp signals rather than their explicit formula. We define graph chirp signals as the inverse GFRFT of a spectral delta function, ensuring that they inherit the core characteristics of chirp signals. Notably, the graph chirp signal defined by Chen et al. involves only the graph chirp rate, without incorporating a notion of graph initial frequency. In contrast, our definition allows for both the graph chirp rate and graph initial frequency, mirroring the dual-parameter structure of classical chirp signals. Furthermore, their graph chirp signal does not satisfy the GFRFT-invariance, and due to the absence of graph initial frequency, it also fails to exhibit orthogonality. The comparison of two kinds of graph chirp signals is illustrated in Fig. 2.
Figure 2: The comparison of two kind of graph chirp signals.
@@
$\mathbf{A}=\left[\begin{array}[]{llll}&&&1\\ 1&&&\\ &\ddots&&\\ &&1&\end{array}\right].$ (21)
On a cycle graph, the GFRFT matrix, $\mathbf{F}_{G}^{a}$, is equal to the discrete fractional Fourier transform (DFRFT) matrix [17]. Consequently, the graph chirp signals on a cycle graph are defined as the columns of the DFRFT matrix with a fractional order $-a$.
The GFRFT of a graph chirp signal $\mathbf{u}_{k}^{a}$ produces a corresponding impulse in the graph fractional spectral domain, analogous to the chirp signal’s behavior in classical SP. For simplicity, let $\mathbf{U}_{a}=(\mathbf{u}_{1}^{a},\mathbf{u}_{2}^{a},\cdots,\mathbf{u}_{N}^{a})$. For example, Fig. 3 illustrates the GFRFT of the chirp signal $\mathbf{u}_{50}^{0.5}$ on the David sensor, with the fractional order $a$ varying from 0.10.10.1 to 111. It can be seen that the GFRFT for this fractional order generates an impulse at the chirp rate $a=0.5$, indicating that the signal exhibits a well-defined frequency concentration at this order. At other fractional orders, the spectrum is more mixed. The GFRFT effectively handles graph chirp signals by providing a flexible method to adjust the order. This aligns with the classical treatment of chirp signals, where the frequency evolution over time is captured precisely.
Figure 3: GFRFTs of the chirp signal $\mathbf{u}_{50}^{0.5}$ on David sensor at different fractional orders.
IV Fractional vertex–frequency energy distribution
IV-A Vertex–frequency energy distribution
In classical SP, the WD [21, 22] is a well-known tool that provides a joint representation of a signal, offering valuable insight into its time–frequency characteristics. It is particularly effective for the analysis and processing of chirp signals, such as filtering and detection. In GSP, the concept of energy distribution holds an analogous importance to the role of WD in classical SP.
In [25], the GED is proposed, which corresponds to the Rihaczek distribution in time–frequency analysis and is defined as follows: